OpenStudy (anonymous):

Complete the square and use quadratic formula: x^2 - 2x - 1 = 0 I got to: x^2 - 2x = 1 (x + 2)^2 = +/-√2

OpenStudy (bahrom7893):

x^2 - 2x - 1 + 2 = 2 (x-1)^2 = 2

OpenStudy (anonymous):

Darn. (x+2)^2 = 2 x + 2 = +/-√2

OpenStudy (bahrom7893):

why would they want to complete the square? it's not helping..

OpenStudy (bahrom7893):

x-1 = +-sqrt(2) x = -sqrt(2) + 1 x = sqrt(2)+1

OpenStudy (bahrom7893):

well actually it was helping

OpenStudy (bahrom7893):

That's by completing the square

OpenStudy (bahrom7893):

x^2 - 2x - 1 = 0 D = (-2)^2 - 4*1*(-1)=4+4 = 8

OpenStudy (bahrom7893):

x1 = (2 - sqrt(8))/2 = 1 - sqrt(2) x2 = (2 + sqrt(8))/2 = 1 + sqrt(2)

OpenStudy (bahrom7893):

come on now, make me level 59!

OpenStudy (bahrom7893):

yay lol!

OpenStudy (anonymous):

Thank you. :D

OpenStudy (bahrom7893):

tnx

OpenStudy (bahrom7893):

and np